### v1

@@ 1. Introduction
Let $F$ be a field and let $G_{F}$ denote its absolute Galois group. For a prime number $p$, let $\mathcal{C}^{\bullet}(G_{F},\mathbb{F}_{p})$ denote the differential graded algebra of inhomogeneous continuous cochains of $G_{F}$ with coefficients in the constant discrete $G_{F}$-module $\mathbb{F}_{p}$. In [13], Hopkins and Wickelgren showed that all triple Massey products of local and global fields at the prime $2$ vanish whenever they are defined. Since triple Massey products are the first obstruction to formality, Hopkins–Wickelgren therefore asked in [13, Question 1.4] whether $\mathcal{C}^{\bullet}(G_{F},\mathbb{F}_{2})$ is formal, i.e., whether there is a zigzag of quasi-isomorphisms of differential graded algebras between $\mathcal{C}^{\bullet}(G_{F},\mathbb{F}_{2})$ and its cohomology $H^{\bullet}(G_{F},\mathbb{F}_{2})$. However, Positselski showed in [37, Section 9.11] and [38, §6] that $\mathcal{C}^{\bullet}(G_{F},\mathbb{F}_{2})$ is not formal in general for local fields. Then Harpaz–Wittenberg in [11, Example A.15] and more recently Merkurjev–Scavia in [26, Theorem 1.6] and [28, Theorem 1.3] provided examples which show that the second obstruction to formality is not trivial in general, i.e., not all fourfold Massey products are defined when the neighbouring cup-products vanish.
@@
where $x,y]=x^{-1}y^{-1}xy$ denotes the commutator of elements $x,y\in G$. Our main result is the following theorem, see Theorems 6.6 and 6.12:
@@
Demushkin groups are Poincaré groups of dimension two and play an important role for example in number theory since the maximal pro-$p$ quotients of absolute Galois groups of local fields that contain a primitive $p$-th root of unity are Demushkin groups or are trivial (see for example [41, Théorème 4.2]). Demushkin groups are fundamental building blocks of the class of elementary type pro-$p$ groups in the sense of Efrat. These groups are defined inductively as the class of pro-$p$ groups that includes finitely generated free pro-$p$ groups and Demushkin groups and is closed under taking free pro-$p$ products and semi-direct products with $\mathbb{Z}_{p}$. For the latter one requires that all groups are equipped with an orientation (see [6, Section 3] and e.g. [32, Section 4]). For positive results on Efrat’s elementary type conjecture see [7], [8], [10, Chapter 12]. Demushkin groups also arise as pro-$p$ completions of fundamental groups of compact surfaces $\Sigma$ of genus $g\geq 1$ when $\Sigma$ is orientable and $g\geq 2$ when $\Sigma$ is not orientable (see also [21]).
Before we outline the proof of Theorem 1.2, we describe the relation of our work to the Massey vanishing conjecture. Mináč and Tân conjectured in [29, Conjecture 1.1] that, for every field $F$ and prime $p$, $G_{F}$ satisfies $n$-Massey vanishing with respect to $p$, i.e., all $n$-fold Massey products of elements in $H^{1}(G_{F},\mathbb{F}_{p})$ vanish whenever they are defined. By the work of Matzri [24], Efrat–Matzri [9] and Mináč–Tân [30], all fields satisfy triple Massey vanishing with respect to all primes. In [12], Harpaz–Wittenberg showed that number fields satisfy $n$-Massey vanishing with respect to all primes. More recently, Merkurjev–Scavia proved in [27] that all fields satisfy fourfold Massey vanishing with respect to $p=2$. Other cases of the conjecture have been proven in [34], [35] and [39]. The vanishing of Massey products has concrete consequences for the structure of the Zassenhaus filtration of an absolute Galois and thereby led to new examples of profinite groups which are not absolute Galois groups of a field (see for example [30, 31]).
We say that a profinite group $G$ satisfies strong $n$-Massey vanishing with respect to $p$ if all $n$-fold Massey products of elements in $H^{1}(G,\mathbb{F}_{p})$ are defined and vanish when the neighbouring cup-products vanish. For $n\geq 4$, this is a strictly stronger condition. By the work of Pál–Szabó [35, Theorem 3.5] (see also [23, Theorem 3.1 and its proof]), pro-$p$ Demushkin groups satisfy strong $n$-Massey vanishing with respect to $p$ and all $n\geq 3$. Moreover, by [23, Theorem 1], the absolute Galois groups of number fields which do not contain a primitive $p$th root of unity satisfy strong $n$-Massey vanishing with respect to $p$ for all $n\geq 3$. Strong vanishing of triple Massey products is a necessary condition for the $A_{3}$-formality of $\mathcal{C}^{\bullet}(G,\mathbb{F}_{p})$. We note, however, that $A_{3}$-formality is a significant strengthening of the vanishing of the Massey product obstructions since $A_{3}$-formality requires that a specific element in the triple Massey product of elements in $H^{\bullet}(G,\mathbb{F}_{p})$ vanishes and that defining systems of triple Massey products can be chosen compatibly.
@@
Let $p=3$ and let $G$ be the pro-$p$ group with generators $x_{1},x_{2},x_{3},x_{4}$ subject to the single relation $1=x_{1}^{3}[x_{1},x_{2}][x_{3},x_{4}]$. Following [6] and [19, page 254], the group $G$ is realisable as the maximal pro-$3$ Galois group $G_{F}(3)$ of the field $F=\mathbb{Q}_{3}(\zeta_{3})$ where $\zeta_{3}$ is a root of unity of order $3$. Thus, Theorem 1.2 shows that $\mathcal{C}^{\bullet}(G_{F},\mathbb{F}_{p})$ is not $A_{3}$-formal for $F=\mathbb{Q}_{3}(\zeta_{3})$.
@@
For $p\geq 5$, are all elementary type pro-$p$ groups $A_{3}$-formal?
We note that free pro-$p$ groups are $A_{\infty}$-formal (see Proposition 5.3). We can also show that the free pro-$p$ product of pro-$p$ groups which are Koszul and $A_{3}$-formal is again $A_{3}$-formal. Moreover, by [32, Theorem A], the $\mathbb{F}_{p}$-cohomology algebra of elementary type pro-$p$ groups is Koszul. Hence, the only case missing for an answer to Question 1.6 is the one of semi-direct products.
Based on the results of [23], we also ask the following
@@
The proof of Theorem 1.2 relies on the fact that the quadratic algebra $H^{\bullet}(G,\mathbb{F}_{p})$ for a pro-$p$ Demushkin group is Koszul. The latter is known by the work of Mináč–Pasini–Quadrelli–Tân in [32, Theorem 5.2]. In fact, they prove the stronger result that $H^{\bullet}(G,\mathbb{F}_{p})$ is a PBW-algebra which implies that $H^{\bullet}(G,\mathbb{F}_{p})$ is Koszul. We provide an alternative proof that $H^{\bullet}(G,\mathbb{F}_{p})$ is Koszul in Section 5.3 for completeness and convenience of the reader.
@@
In the proof of Theorem 1.2 we make frequent use of the assumption that $p$ is odd. We do not know whether pro-$2$ Demushkin groups are $A_{3}$-formal or not. For $q\geq 5$, we do not know for which integer $n>3$ pro-$p$ Demushkin groups are not $A_{n}$-formal with respect to $p$. That such an $n$ exists follows from Positselski’s work in [38, §6].
@@ Acknowledgements
The second-named author would like to thank Mads Hustad Sandøy for helpful conversations on $A_{\infty}$-algebras.
@@ 2.1. $A_{3}$-algebras
This is a well-known result for $A_{\infty}$-algebras (see for example [44, Proof of Theorem 7.2.2]). We provide a sketch of the proof since it provides us with constructions that will be used later. Moreover, we could not find a proof of the uniqueness statement for $A_{3}$-algebras in the literature.
Proof of Theorem 2.7.
@@ 2.3. $A_{3}$-formality and Massey products
We note that by [14, Theorem 3.4] it is not true that for the fourfold Massey product $\langle a_{1},a_{2},a_{3},a_{4}\rangle$ to be defined it is not sufficient that the triple Massey products $\langle a_{1},a_{2},a_{3}\rangle$ and $\langle a_{2},a_{3},a_{4}\rangle$ vanish. There is no obstruction, however, for an $A_{3}$-formal differential graded algebra as the following result shows.
@@
Similar to [14, Section 4], we can construct a differential graded algebra for which all triple Massey products vanish, but which is not $A_{3}$-formal, as follows. Let $S$ be the set $\{a_{12},a_{23},a_{34},a_{45},a_{13},a_{24},a^{\prime}_{24},a_{35},a_{14},a^{\prime}_{25}\}$, and let $V$ be the free $\mathbb{F}$-vector space on the set $S$. Let $\mathcal{A}$ be the differential graded algebra over $\mathbb{F}$ whose underlying algebra is the graded tensor algebra $T(V)$ with $V$ in degree one and differential the unique derivation defined on generators as listed in the following table:
@@ 3.2. Kadeishvili’s criterion for $A_{3}$-formality
The following result is a modified version of Kadeishvili’s theorem [16] (see also [40, Theorem 4.7]).
@@
Let $H^{\bullet}$ be a graded algebra with $\mathrm{HH}^{3,-1}(H^{\bullet})=0$. Then Theorem 3.9 implies that $H^{\bullet}$ is intrinsically $A_{3}$-formal, i.e., every connected differential graded algebra $\mathcal{A}$ with cohomology isomorphic to $H^{\bullet}$ is $A_{3}$-formal. See also [34, Lemma 4.6].
@@ 4.1. Koszul algebras
Let $(B(A),d_{B})$ denote the bar complex of $A$ (see e.g. [44, (1.1.4)]). Since $R$ describes the relations in $A$, it follows that the natural inclusion $K(A)\hookrightarrow B(A)$ defines a morphism of complexes.
@@
We note that there are many different ways to characterise the notion of a Koszul algebra. We refer to [36, Chapter 2, Sections 1 and 3] and for example [1, Section 2] and [43, Section 3] for proofs that the alternative definitions are equivalent to Definition 4.3.
@@
Important examples of Koszul algebras include symmetric and exterior algebras over $\mathbb{F}$ (see [36, Example on page 20]).
@@ 4.2. Hochschild cohomology of Koszul algebras
We then get the following result (see also [43, Proposition 3.3]):
@@
Let $A^{e}=A\otimes A^{\mathrm{op}}$ denote the enveloping algebra of $A$ where $A^{\mathrm{op}}$ denotes the opposite algebra of $A$ (see e.g., [44, Section 1.1]). The ring $A^{e}$ inherits a structure of a graded $\mathbb{F}$-algebra from $A$. We can consider the graded $A$-bimodule $M$ as a graded $A^{e}$-module. For every $n$, there is a natural isomorphism
@@
It is well-known that $A$ is a Koszul algebra (see e.g. [36, Example on page 20]). We have $K_{2}^{2}(A)=R$, and
@@
Let $A$ again be the exterior algebra on two generators $x$ and $y$ in degree one. By [22, Proposition 3.2], in order to define an $A_{\infty}$-algebra structure on the graded algebra $A$, it suffices to specify $\mathbb{F}$-linear maps $m_{n}\colon(A^{1})^{\otimes n}\to A^{2}$ for all $n\neq 2$. We set $m_{1}=0$, $m_{n}=0$ for $n\geq 4$ and can choose $m_{3}$ to be any $\mathbb{F}$-linear map $(A^{1})^{\otimes 3}\to A^{2}$. This defines a minimal $A_{\infty}$-algebra structure on $A$. Now we recall that, for every $A_{\infty}$-algebra $\mathcal{A}$, the canonical morphism of $A_{\infty}$-algebras $\mathcal{A}\to\Omega B\mathcal{A}$, where $\Omega$ denotes the cobar and $B$ the bar construction, is a quasi-isomorphism (see for example [22, Proposition 4.5]). Since $\Omega B\mathcal{A}$ is a differential graded algebra, every $A_{\infty}$-algebra is quasi-isomorphic as an $A_{\infty}$-algebra to a differential graded algebra. Since quasi-isomorphisms between $A_{\infty}$-algebras are homotopy equivalences by [18, Theorem 3.7 on page 13], we conclude that $A$ is the minimal model of some differential graded algebra $\mathcal{A}$ with canonical class $m_{3}]\in\mathrm{HH}^{3,-1}(\mathcal{A})$. By Theorem 2.7 and Example 4.7, this shows that there are many differential graded algebras whose cohomology algebra is isomorphic to $A$ but which are not quasi-isomorphic as $A_{3}$-algebras. In particular, $A$ is not intrinsically $A_{3}$-formal.

### v2

@@ 1. Introduction
Let $F$ be a field and let $G_{F}$ denote its absolute Galois group. For a prime number $p$, let $\mathcal{C}^{\bullet}(G_{F},\mathbb{F}_{p})$ denote the differential graded algebra of inhomogeneous continuous cochains of $G_{F}$ with coefficients in the constant discrete $G_{F}$-module $\mathbb{F}_{p}$. In [14], Hopkins and Wickelgren showed that all triple Massey products of local and global fields at the prime $2$ vanish whenever they are defined. Since triple Massey products are the first obstruction to formality, Hopkins–Wickelgren therefore asked in [14, Question 1.4] whether $\mathcal{C}^{\bullet}(G_{F},\mathbb{F}_{2})$ is formal, i.e., whether there is a zigzag of quasi-isomorphisms of differential graded algebras between $\mathcal{C}^{\bullet}(G_{F},\mathbb{F}_{2})$ and its cohomology $H^{\bullet}(G_{F},\mathbb{F}_{2})$. However, Positselski showed in [42, Section 9.11] and [44, §6] that $\mathcal{C}^{\bullet}(G_{F},\mathbb{F}_{2})$ is not formal in general for local fields. Then Harpaz–Wittenberg in [12, Example A.15] and more recently Merkurjev–Scavia in [27, Theorem 1.6] and [29, Theorem 1.3] provided examples which show that the second obstruction to formality is not trivial in general, i.e., not all fourfold Massey products are defined when the neighbouring cup-products vanish.
@@
where $x,y]=x^{-1}y^{-1}xy$ denotes the commutator of elements $x,y\in G$.
Our main result is the following theorem, see Theorems 6.6 and 6.12:
@@
Demushkin groups are Poincaré groups of dimension two and play an important role for example in number theory since the maximal pro-$p$ quotients of absolute Galois groups of local fields that contain a primitive $p$-th root of unity are Demushkin groups or are trivial (see for example [47, Théorème 4.2]). Demushkin groups are fundamental building blocks of the class of elementary type pro-$p$ groups in the sense of Efrat. These groups are defined inductively as the class of pro-$p$ groups that includes finitely generated free pro-$p$ groups and Demushkin groups and is closed under taking free pro-$p$ products and semi-direct products with $\mathbb{Z}_{p}$. For the latter one requires that all groups are equipped with an orientation (see [7, Section 3] and e.g. [31, Section 4]). For positive results on Efrat’s elementary type conjecture see [8], [9], [11, Chapter 12]. Demushkin groups also arise as pro-$p$ completions of fundamental groups of compact surfaces $\Sigma$ of genus $g\geq 1$ when $\Sigma$ is orientable and $g\geq 2$ when $\Sigma$ is not orientable (see also [22]).
@@
Let $p=3$ and let $G$ be the pro-$p$ group with generators $x_{1},x_{2},x_{3},x_{4}$ subject to the single relation $1=x_{1}^{3}[x_{1},x_{2}][x_{3},x_{4}]$. Following [7] and [20, page 254], the group $G$ is realisable as the maximal pro-$3$ Galois group $G_{F}(3)$ of the field $F=\mathbb{Q}_{3}(\zeta_{3})$ where $\zeta_{3}$ is a root of unity of order $3$. Thus, Theorem 1.2 shows that $\mathcal{C}^{\bullet}(G_{F},\mathbb{F}_{p})$ is not $A_{3}$-formal for $F=\mathbb{Q}_{3}(\zeta_{3})$.
@@
Relation to the Massey vanishing conjecture of Mináč–Tân
Before we outline the proof of Theorem 1.2, we describe the relation of our work to the Massey vanishing conjecture. Mináč and Tân conjectured in [33, Conjecture 1.1] that, for every field $F$ and prime $p$, $G_{F}$ satisfies $n$-Massey vanishing with respect to $p$, i.e., all $n$-fold Massey products of elements in $H^{1}(G_{F},\mathbb{F}_{p})$ vanish whenever they are defined. By the work of Matzri [25], Efrat–Matzri [10] and Mináč–Tân [35], all fields satisfy triple Massey vanishing with respect to all primes. In [13], Harpaz–Wittenberg showed that number fields satisfy $n$-Massey vanishing with respect to all primes. More recently, Merkurjev–Scavia proved in [28] that all fields satisfy fourfold Massey vanishing with respect to $p=2$. Other cases of the conjecture have been proven in [39], [40] and [45]. The vanishing of Massey products has concrete consequences for the structure of the Zassenhaus filtration of an absolute Galois and thereby led to new examples of profinite groups which are not absolute Galois groups of a field (see for example [35, 36]).
In [34, Definition 4.5], Mináč and Tân formulate the following related property. Let $G$ be a profinite group and $p$ be a prime number. Then $G$ is said to have the cup-defining $n$-fold Massey product property (with respect to $\mathbb{F}_{p}$) if for every $\chi_{1},\ldots,\chi_{n}\in H^{1}(G,\mathbb{F}_{p})$ with $0=\chi_{1}\cup\chi_{2}=\chi_{2}\cup\chi_{3}=\cdots=\chi_{n-1}\cup\chi_{n}$ the $n$-fold Massey product $\langle\chi_{1},\ldots,\chi_{n}\rangle$ is defined. For $n\geq 4$, this is a non-trivial condition, and, in [34, Remark 4.4], Mináč–Tân show that not all pro-$p$ groups have the cup-defining $n$-fold Massey product property with respect to $\mathbb{F}_{p}$. In [34, Question 4.2], Mináč–Tân ask whether every Galois group of a maximal $p$-extension of a field $F$ containing a primitive $p$-th root of unity has the cup-defining $n$-fold Massey product property with respect to $\mathbb{F}_{p}$ (see also [37, Section 8]). They show that, for two pro-$p$ groups $G_{1}$ and $G_{2}$, the free pro-$p$ product $G_{1}\ast G_{2}$ has the cup-defining $n$-fold Massey product property with respect to $\mathbb{F}_{p}$ if and only if both $G_{1}$ and $G_{2}$ do. Moreover, in [34, Proposition 4.1], Mináč–Tân prove that pro-$p$ Demushkin groups have the cup-defining $n$-fold Massey product property with respect to $\mathbb{F}_{p}$. Together with their work in [36], this implies that pro-$p$ Demushkin groups have the following stronger property.
We say that a profinite group $G$ satisfies strong $n$-Massey vanishing with respect to $p$ if for every $\chi_{1},\ldots,\chi_{n}\in H^{1}(G,\mathbb{F}_{p})$ with $0=\chi_{1}\cup\chi_{2}=\cdots=\chi_{n-1}\cup\chi_{n}$ the $n$-fold Massey product $\langle\chi_{1},\ldots,\chi_{n}\rangle$ is defined and vanishes. For $n\geq 4$, this is a strictly stronger condition than $n$-Massey vanishing. By the work of Mináč–Tân in [34, Proposition 4.1] and [35, Theorem 4.3], pro-$p$ Demushkin groups satisfy strong $n$-Massey vanishing with respect to $p$ and all $n\geq 3$. An independent proof that pro-$p$ Demushkin groups satisfy strong $n$-Massey vanishing for all $n\geq 3$ was given by Pál–Szabó in [40, Theorem 3.5]. Moreover, by [24, Theorem 1], the absolute Galois groups of number fields which do not contain a primitive $p$th root of unity satisfy strong $n$-Massey vanishing with respect to $p$ for all $n\geq 3$. Strong vanishing of triple Massey products is a necessary condition for the $A_{3}$-formality of $\mathcal{C}^{\bullet}(G,\mathbb{F}_{p})$. We note, however, that $A_{3}$-formality is a significant strengthening of the vanishing of the Massey product obstructions since $A_{3}$-formality requires that a specific element in the triple Massey product of elements in $H^{\bullet}(G,\mathbb{F}_{p})$ vanishes and that defining systems of triple Massey products can be chosen compatibly.
New questions
@@
Let $p\geq 5$ be a prime number. Are all elementary type pro-$p$ groups $A_{3}$-formal?
We note that free pro-$p$ groups are $A_{\infty}$-formal (see Proposition 5.3). We can also show that the free pro-$p$ product of pro-$p$ groups which are Koszul and $A_{3}$-formal is again $A_{3}$-formal. Moreover, by [31, Theorem A], the $\mathbb{F}_{p}$-cohomology algebra of elementary type pro-$p$ groups is Koszul. Hence, the only case missing for an answer to Question 1.6 is the one of semi-direct products. Based on the results of [24], we also ask the following
@@
The proof of Theorem 1.2 relies on the fact that the quadratic algebra $H^{\bullet}(G,\mathbb{F}_{p})$ for a pro-$p$ Demushkin group is Koszul. The latter is known by the work of Mináč–Pasini–Quadrelli–Tân in [31, Theorem 5.2]. In fact, they prove the stronger result that $H^{\bullet}(G,\mathbb{F}_{p})$ is a PBW-algebra which implies that $H^{\bullet}(G,\mathbb{F}_{p})$ is Koszul. We provide an alternative proof that $H^{\bullet}(G,\mathbb{F}_{p})$ is Koszul in Section 5.3 for completeness and convenience of the reader. We recall that Positselski and Voevodsky conjectured that all Galois cohomology algebras are Koszul [43, §0.1, page 128]. Further results on the Koszulity of Galois cohomology algebras we refer to [30], [31], [32], [39], and [43].
@@
In the proof of Theorem 1.2 we make frequent use of the assumption that $p$ is odd. We do not know whether pro-$2$ Demushkin groups are $A_{3}$-formal or not. For $q\geq 5$, we do not know for which integer $n>3$ pro-$p$ Demushkin groups are not $A_{n}$-formal with respect to $p$. That such an $n$ exists follows from Positselski’s work in [44, §6].
@@ Acknowledgements
We are very grateful to Jan Mináč for helpful comments. GQ would also like to thank Mads Hustad Sandøy for valuable conversations on $A_{\infty}$-algebras.
@@ 2.1. $A_{3}$-algebras
This is a well-known result for $A_{\infty}$-algebras (see for example [50, Proof of Theorem 7.2.2]). We provide a sketch of the proof since it provides us with constructions that will be used later. Moreover, we could not find a proof of the uniqueness statement for $A_{3}$-algebras in the literature.
Proof of Theorem 2.7..
@@ 2.3. $A_{3}$-formality and Massey products
We note that by [15, Theorem 3.4] it is not true that for the fourfold Massey product $\langle a_{1},a_{2},a_{3},a_{4}\rangle$ to be defined it is not sufficient that the triple Massey products $\langle a_{1},a_{2},a_{3}\rangle$ and $\langle a_{2},a_{3},a_{4}\rangle$ vanish. There is no obstruction, however, for an $A_{3}$-formal differential graded algebra as the following result shows.
@@
Similar to [15, Section 4], we can construct a differential graded algebra for which all triple Massey products vanish, but which is not $A_{3}$-formal, as follows. Let $S$ be the set $\{a_{12},a_{23},a_{34},a_{45},a_{13},a_{24},a^{\prime}_{24},a_{35},a_{14},a^{\prime}_{25}\}$, and let $V$ be the free $\mathbb{F}$-vector space on the set $S$. Let $\mathcal{A}$ be the differential graded algebra over $\mathbb{F}$ whose underlying algebra is the graded tensor algebra $T(V)$ with $V$ in degree one and differential the unique derivation defined on generators as listed in the following table:
@@ 3.2. Kadeishvili’s criterion for $A_{3}$-formality
The following result is a modified version of Kadeishvili’s theorem [17] (see also [46, Theorem 4.7]).
@@
Let $H^{\bullet}$ be a graded algebra with $\mathrm{HH}^{3,-1}(H^{\bullet})=0$. Then Theorem 3.9 implies that $H^{\bullet}$ is intrinsically $A_{3}$-formal, i.e., every connected differential graded algebra $\mathcal{A}$ with cohomology isomorphic to $H^{\bullet}$ is $A_{3}$-formal. See also [39, Lemma 4.6]. We refer to [5] for a non-trivial example of an intrinsically $A_{3}$-formal $\mathbb{F}_{2}$-algebra.
@@ 4.1. Koszul algebras
Let $(B(A),d_{B})$ denote the bar complex of $A$ (see e.g. [50, (1.1.4)]). Since $R$ describes the relations in $A$, it follows that the natural inclusion $K(A)\hookrightarrow B(A)$ defines a morphism of complexes.
@@
We note that there are many different ways to characterise the notion of a Koszul algebra. We refer to [41, Chapter 2, Sections 1 and 3] and for example [1, Section 2] and [49, Section 3] for proofs that the alternative definitions are equivalent to Definition 4.3.
@@
Important examples of Koszul algebras include symmetric and exterior algebras over $\mathbb{F}$ (see [41, Example on page 20]).
@@ 4.2. Hochschild cohomology of Koszul algebras
We then get the following result (see also [49, Proposition 3.3]):
@@
Let $A^{e}=A\otimes A^{\mathrm{op}}$ denote the enveloping algebra of $A$ where $A^{\mathrm{op}}$ denotes the opposite algebra of $A$ (see e.g., [50, Section 1.1]). The ring $A^{e}$ inherits a structure of a graded $\mathbb{F}$-algebra from $A$. We can consider the graded $A$-bimodule $M$ as a graded $A^{e}$-module. For every $n$, there is a natural isomorphism
@@
It is well-known that $A$ is a Koszul algebra (see e.g. [41, Example on page 20]). We have $K_{2}^{2}(A)=R$, and
@@
Let $A$ again be the exterior algebra on two generators $x$ and $y$ in degree one. By [23, Proposition 3.2], in order to define an $A_{\infty}$-algebra structure on the graded algebra $A$, it suffices to specify $\mathbb{F}$-linear maps $m_{n}\colon(A^{1})^{\otimes n}\to A^{2}$ for all $n\neq 2$. We set $m_{1}=0$, $m_{n}=0$ for $n\geq 4$ and can choose $m_{3}$ to be any $\mathbb{F}$-linear map $(A^{1})^{\otimes 3}\to A^{2}$. This defines a minimal $A_{\infty}$-algebra structure on $A$. Now we recall that, for every $A_{\infty}$-algebra $\mathcal{A}$, the canonical morphism of $A_{\infty}$-algebras $\mathcal{A}\to\Omega B\mathcal{A}$, where $\Omega$ denotes the cobar and $B$ the bar construction, is a quasi-isomorphism (see for example [23, Proposition 4.5]). Since $\Omega B\mathcal{A}$ is a differential graded algebra, every $A_{\infty}$-algebra is quasi-isomorphic as an $A_{\infty}$-algebra to a differential graded algebra. Since quasi-isomorphisms between $A_{\infty}$-algebras are homotopy equivalences by [19, Theorem 3.7 on page 13], we conclude that $A$ is the minimal model of some differential graded algebra $\mathcal{A}$ with canonical class $m_{3}]\in\mathrm{HH}^{3,-1}(\mathcal{A})$. By Theorem 2.7 and Example 4.7, this shows that there are many differential graded algebras whose cohomology algebra is isomorphic to $A$ but which are not quasi-isomorphic as $A_{3}$-algebras. In particular, $A$ is not intrinsically $A_{3}$-formal (see Remark 3.12).